OpenStudy (anonymous):

What formula is used to find the theoretical probability of an event? P(event) =

OpenStudy (nurali):

The probability of an event is the number of ways the event can occur divided by the total number of possible outcomes. For example, when you toss a die, there are 6 possible outcomes. Suppose that the event is tossing a 5 or greater. There are two ways that event can occur (rolling a 5 or a 6), so the probability is 2/6 = 1/3.

OpenStudy (anonymous):

Okay.. Does that work with 12 cards also? @Nurali

OpenStudy (nurali):

If you had 12 different possible outcomes and they each had an equal chance of happening (eg a box with 12 cards inside and a person picks out one of them) then the probability of, for example, picking a red card would be the number of red cards divided by the total number of possible outcomes. So if there were 2 red cards out of the 12, p(picking a red card) = 2/12 = 1/6.